I need to find the equation of the line containg (-2, -7) and (-5, -8)
Answer by mananth(16946)

x1 y1 x2 y2
-2 -7 -5 -8

slope m =(y2-y1)/(x2-x1) 0.33
(-8-(-7 ))/(-5-(-2))
(-1/-3)
m=0.33

Plug value of the slope and point (-2,-7) in
Y = m x + b
-7.00 = -0.67 + b
b= -7-0.67
b= -6.33
So the equation will be
Y =0.33 x-6.33